OpenStudy (pokiedokie):

please help just a quick question i want to be sure on!! f(x) = -x + 4

OpenStudy (jdoe0001):

question? we kinda missed that part nice equation though

OpenStudy (pokiedokie):

okay so for each function, find f(-2), f(-0,5), and f(3) i'm only confused because i'm not sure if the value would work with x f(-2)= -x + 4 or if it would work with the f function (-2)(x) = -x + 4 im know im making it more confusing for myself but i just need to be validated on which one

OpenStudy (pokiedokie):

i know im supposed to distribute but that's it

OpenStudy (pokiedokie):

if i'm being confusing just tell me and I can elaborate more, sorry

OpenStudy (jdoe0001):

\(\begin{array}{cllll} f(x)=-x+4 \\\hline\\ f({\color{brown}{ -2}})=-({\color{brown}{ -2}})+4\\ f({\color{brown}{ -0.5}})=-({\color{brown}{ -0.5}})+4\\ f({\color{brown}{ -3}})=-({\color{brown}{ -3}})+4 \end{array}\)

OpenStudy (pokiedokie):

ahhh thank you!! i knew i was just making it harder than it actually was, thank you so much!!

OpenStudy (jdoe0001):

yw